OpenStudy (anonymous):

For the following system, if you isolated x in the first equation to use the Substitution Method, what expression would you substitute into the second equation? -x - 2y = -4 3x + y = 12 A. -2y - 4 B. 2y - 4 C. 2y + 4 D. -2y + 4

OpenStudy (ankit042):

do you know how to proceed?

OpenStudy (anonymous):

No not really

OpenStudy (anonymous):

But i would choose A

OpenStudy (ankit042):

how did you reach A can you please explain?

OpenStudy (anonymous):

-x - 2y = -4 Equation 1 3x + y = 12 Equation 2 To isolate x in Equation 1, you need to add 2y to both sides of the equal sign. -x - 2y + 2y = -4 + 2y -x = -4 + 2y And that is equation 3 Before we can use the equation 3 substitution, we have a slight problem to correct. Notice that equation 3 is in terms of negative x, and equation 2 has a positive x terrm. So to change the polarity of the x term in equation 3, multiply both sides by (-1) (-1)(-x) = (-1)(-4 + 2y) x = 4 - 2y Now just substitute (4 - 2y) for x in equation 2. What do you get @shannon342

OpenStudy (anonymous):

-2y+4?

OpenStudy (anonymous):

since x is negative in the first equation, should the second be also negative?

OpenStudy (anonymous):

Yes, that is what I got

OpenStudy (anonymous):

Thanks for the help :)

OpenStudy (anonymous):

np can i get a medal :) ?

OpenStudy (anonymous):

ty